OpenStudy (lgbasallote):

\[\mathcal L^{-1} \left \{\frac{2s^2 + 1}{(s+1)^2}\right\} = 1 + e^{-t} - 3te^{-t}?\]

OpenStudy (anonymous):

take laplace of right hand side and check .are you are getting the original Question.?

OpenStudy (lgbasallote):

\[\frac{1}{s} + \frac{1}{s-1} -\frac{3}{(s-1)^2}??\]

OpenStudy (anonymous):

it means something is wrong :P you are not getting original question. are you?

OpenStudy (lgbasallote):

well that's what i got when i turned it into partial fractions

OpenStudy (anonymous):

Before Partial fraction make sure it is proper fraction ?

OpenStudy (lgbasallote):

sounds proper to me

OpenStudy (lgbasallote):

uhmm wait...that should be s+1 not s-1

OpenStudy (anonymous):

if not long division . i do not think it is proper. long division is required. Because the degree of both numerator and denominator polynomial is same ??

OpenStudy (lgbasallote):

...well i still think it will come out same :/

OpenStudy (anonymous):

wolfram also admits it is improper and do long division first. http://www.wolframalpha.com/input/?i=Partial+fractions++%282s%5E2%2B1%29%2F%28s%2B1%29%5E2%29

OpenStudy (lgbasallote):

seems im wrong >.<

OpenStudy (anonymous):

yes it seems :(

OpenStudy (anonymous):

so you should first perform long division. right ?

OpenStudy (lgbasallote):

ugh i dont know anymore =_= i think i'll retire for now....this is the second to the last topic anyway @_@

OpenStudy (anonymous):

if you want . but an interesting function is going to come out of this inverse Laplace; wanna know ?

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

ok have you ever heard of inverse Laplace of 1??. it is here \[\Large L^{-1}(1)=\delta(t)\] where\[\Large \delta(t)\] it is delta function. when you will perform long division and partial fraction you will have. \[\Large +2-\frac{4}{s+1}+\frac{3}{(s+1)^2}\] i hope you can find the inverse of second and third term for the first term you will have \[\Large 2L^{-1}(1)=2\delta(t)\]

OpenStudy (lgbasallote):

i have no idea what you just said..

OpenStudy (lgbasallote):

but since it's \[\frac{-4}{s+1} + \frac{3}{(s+1)^2} +2\] i assume inverse laplace would be \[-4e^{-t} + 3te^{-t} + \frac2s\] correct?

OpenStudy (anonymous):

the first two terms are correct but third one is not. how an s can be there when you have moved to the t domain ? always remember when you take Laplace all t should be converted to s and when you take inverse all s should be back into t.

OpenStudy (lgbasallote):

hmm..im pretty sure it's 2/s though :/

OpenStudy (anonymous):

ok lets go from start!!!! what is Laplace of 1 ?

OpenStudy (lgbasallote):

wolfram saidd it was \[2\delta t\] o.O wth does that even mean

OpenStudy (anonymous):

it mean i was right :)

OpenStudy (lgbasallote):

i think i got meself mixed up..lpalace of 2 is 2/s but it's not the inverse

OpenStudy (anonymous):

good .

OpenStudy (anonymous):

here we goes to delta function! inverse Laplace of 1 is delta function.as i pointed out earlier.

OpenStudy (lgbasallote):

always?

OpenStudy (anonymous):

always !!!!!!

OpenStudy (lgbasallote):

thanks. i learn something new every hour today

OpenStudy (lgbasallote):

too bad i'll forget them in the morning haha

OpenStudy (anonymous):

just one thing more i know you will need this in your questions. what is the Laplace of t ?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

1/s^2 i suppose

OpenStudy (anonymous):

good . since it is transformation i wanted to tell you. if you can move from t to s how to come back from s to t what is inverse Laplace of s? \[\Large L^{-1}(s)=??\]

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

ok just m going to answer it. the inverse Laplace of s is derivative of delta function. \[\Large L^{-1}(s)=\delta'(t)\] i just wanted to tell you this because you will need this in very very near future (may be in your right next question) . so that if you encounter it you will be able to do it !!!:)